# Reasoning Large Language Model Errors Arise from Hallucinating Critical Problem Features

### v1

@@ Abstract
Large language models have recently made great strides in reasoning task performance through chain-of-thought (CoT) strategies trained via reinforcement learning; however, these “reasoning large language models" (RLLMs) remain imperfect reasoners, and understanding the frequencies and causes of their failure modes is important for both users and developers. We test o1-mini, o3-mini, DeepSeek-R1, Claude 3.7 Sonnet, Gemini 2.5 Pro Preview, and Grok 3 Mini Beta on graph coloring as a variable-complexity constraint-satisfaction logic problem, and find evidence from both error rate comparisons and CoT/explanation text analysis that RLLMs are prone to hallucinate edges not specified in the prompt’s description of the graph. This phenomenon persists across multiple problem complexity levels and semantic frames, and it appears to account for a significant fraction of the incorrect answers from every tested model, and the vast majority of them for some models. Our results indicate that RLLMs may possess broader issues with misrepresentation of problem specifics, and we offer suggestions for design choices to mitigate this weakness.
@@ 1 Introduction
Little is publicly known about the precise nature, capabilities, and limitations of RLLM reasoning, due to a combination of RLLMs’ novelty, their rapidly evolving performance frontier, and the frequency with which developers keep model weights and/or training procedure details secret (Zhang et al., 2025); some model APIs even leave the chain of thought completely internal and only report the model’s final answer, likely to prevent competitors from using the chains of thought as training data for their own models. Some research so far has found that despite RLLMs’ improvements over traditional LLMs, they are still not entirely free of flaws such as hallucinating facts in their output, (OpenAI, 2025), reduced performance on problems not found in the training corpus (Mirzadeh et al., 2024), reduced performance when superficially relevant distractor information is added to the prompt (Mirzadeh et al., 2024), imperfect performance even on problems that are simple and algorithmically solvable without fail (Valmeekam et al., 2024; Heyman and Zylberberg, 2025), and decreasing performance with increasing problem complexity even as perfect accuracy remains algorithmically possible (Lin et al., 2025; Heyman and Zylberberg, 2025). Furthermore, an RLLM’s chain of thought may not always accurately report the process by which it arrives at its final answer (Chen et al., 2025b). Gathering more and better evidence about the frequencies and causes of RLLM reasoning failure modes is important both for the growing population of users deploying existing models and for developers seeking to build more capable and reliable models.
@@
Figure 1: Left: A simple undirected graph with 4 vertices and 4 edges. Center: A valid 2-coloring of the graph. Right: A different 4-vertex, 4-edge graph that cannot be 2-colored. Credit to Heyman and Zylberberg (2025).
@@
Some prior work (Stechly et al., 2024; Heyman and Zylberberg, 2025) has evaluated LLM reasoning via graph coloring problems, in which the model must attempt to assign one of $k$ colors to each of the $n$ vertices of a specified undirected graph such that no two vertices with an edge between them receive the same color. (See Figure 1 for a simple example.) What colorings (if any) are valid for a given graph depends on the graph’s precise arrangement of edges. A valid coloring (or proof that there is none) for a given graph and number of available colors $k$ can be found algorithmically without fail, though for $k\geq 3$ the problem is NP-complete. Heyman and Zylberberg (2025) in particular take interest in graph coloring as a problem type that tests both step-by-step reasoning and systematic possibility space exploration, and for which problem instances of widely varying complexity can be easily procedurally generated.
@@
While Stechly et al. (2024) test only the traditional LLM GPT-4 on graph coloring (as well as two other problem types), Heyman and Zylberberg (2025) test the RLLMs o1-mini and DeepSeek-R1 (DeepSeek-AI, 2025a) alongside GPT-4o and three other traditional LLMs. Heyman and Zylberberg (2025) find that while the RLLMs perform much better than the traditional LLMs, the RLLMs remain imperfect even on simple 4-vertex 2-coloring problems, and their error rates increase dramatically as problem complexity increases from 4-vertex 2-coloring to 8-vertex 4-coloring (the highest level they test). In particular, at the 8-vertex level, both RLLMs’ errors mostly consist of claiming that graphs are uncolorable under the constraints when in reality valid colorings exist (as opposed to providing colorings that are in fact invalid, either on the dataset’s uncolorable problems or on its colorable ones). Heyman and Zylberberg (2025) also find a strong positive correlation between error rate and a measure of the difficulty of narrowing down a valid solution to a colorable problem, taking this as evidence that possibility space exploration is RLLMs’ primary area of weakness in the context of problems like graph coloring. However, our own validation experiments have yielded evidence that this correlation is not causally significant (see Appendix A for details), and thus the true primary causes of RLLM errors likely lie elsewhere. Furthermore, the paper does not test any RLLMs released more recently than o1-mini and DeepSeek-R1 due to its publication date, leaving it unclear to what extent its findings generalize to newer RLLMs with generally better benchmark performance.
@@
In this work, we investigate the failure modes of RLLMs on graph coloring problems using not only comparison of error rates and types in the models’ final answers, but also analysis of the justifications they provide for those answers and of their chains of thought when available. Our experiments include o1-mini and DeepSeek-R1 as well as several RLLMs newer and overall better-performing than them. Remarkably, we find that all of the RLLMs we test are prone to hallucinate edges in the graph to be colored that are not listed in the prompt, which can lead them to mistake a colorable graph for an uncolorable one. This phenomenon persists across multiple problem complexity levels and multiple semantic frames in which graph coloring problems can be presented, and it appears to account for at least a significant fraction of the incorrect answers given by every model, and the vast majority of incorrect answers for some models. We present our findings across a series of experiments and conclude by discussing our results’ broader implications for RLLM reasoning and how its flaws might be addressed.
@@ 2 General Methodology
We conduct our experiments on the following RLLMs: (1) OpenAI’s o1-mini (OpenAI, 2024b), specifically o1-mini-2024-09-12, through the OpenAI API; (2) OpenAI’s o3-mini (OpenAI, 2024b), specifically o3-mini-2025-01-31, through the OpenAI API; (3) DeepSeek-R1 (DeepSeek-AI, 2025a) through a combination of the Fireworks AI and DeepInfra online hosts; (4) Anthropic’s Claude 3.7 Sonnet (Anthropic, 2025), specifically claude-3-7-sonnet-20250219 with thinking mode turned on, through the Anthropic API; (5) Google’s Gemini 2.5 Pro Preview Kavukcuoglu (2025) through the Gemini Developer API; and (6) xAI’s Grok 3 Mini Beta xAI (2025) through the xAI API. For o3-mini and Grok 3 Mini Beta, we test all reasoning effort settings (low, medium, and high for the former, and low and high for the latter). We use the mini counterparts of o1, o3, and Grok 3 Beta rather than the full-size models for cost-effectiveness reasons and because the mini models are specialized for logical and mathematical reasoning; we do not expect the broader knowledge of the full-size models to be particularly helpful for the graph coloring task.
For models with variable temperature settings, we set temperature to 0 to match the low-creativity deductive nature of the graph coloring task, with the exception of DeepSeek-R1; its developers recommend a temperature no lower than 0.5 (DeepSeek-AI, 2025b), and we use a temperature of 0.5 for this reason and for consistency with the experimental conditions of Heyman and Zylberberg (2025). We find non-terminating output from our prompts at these temperatures to be infrequent for DeepSeek-R1 and nonexistent for all other models.
@@
We use the same prompting format as Heyman and Zylberberg (2025); the model is given a description of a graph, including a fully enumerated and clearly labeled set of edges, then asked if it is possible to color the graph with a specified set of colors so that no two adjacent vertices receive the same color, and to give a plan for it if so. To broaden our observations, we deploy prompts separately with two different frames, “Math" and “Friends", that Heyman and Zylberberg (2025) designed to probe LLM semantic biases. “Math" prompts present the problem explicitly as coloring a graph with edges connecting numbered vertices; meanwhile, in deliberately less-orthodox “Friends" prompts, the model is given a list of friendships that exist between $n$ people (named alphabetically as Alice, Bob, Carol, etc.) and told that everyone is going to wear a shirt with one of $k$ colors and no one wants to wear the same color shirt as any of their friends. (Examples of the prompts we use can be found in Appendix B.) When evaluating model responses for answer correctness, any responses in which our automatic parser does not detect a coherent answer are reviewed and interpreted manually.
@@
Code for reproducing our experiments and records of our problems, prompts, and responses can be found at https://github.com/AlexHeyman/RLLMGraphColoring. The vast majority of the compute used in our experiments was on the model API side, making it difficult to estimate in terms of total operations. However, each experiment section specifies the number of prompt submissions to the RLLMs involved; the totals across all experiments are 5,155 for each of o1-mini and DeepSeek-R1, 7,651 for Claude 3.7 S, 12,230 for each of o3-mini (high) and Grok 3 MB (high), and 7,230 for each of o3-mini (low), o3-mini (medium), Gemini 2.5 PP, and Grok 3 MB (low). All of our data analysis takes no more than 1 hour to run on a single AMD Ryzen 9 4900HS CPU (maximum FLOPS around $6.4\times 10^{10}$), so as a loose upper bound we estimate that the cost of running the analysis is no more than $2.304\times 10^{14}$ operations.
@@ 3.1 Purpose
Heyman and Zylberberg (2025) find several basic patterns in the errors made by both o1-mini and DeepSeek-R1, including (1) errors on 4-vertex 2-coloring problems are very rare but not impossible; (2) on 8-vertex 4-coloring problems, errors are much more frequent and mostly consist of declaring colorable problems to be uncolorable; and (3) error rates are generally higher for the Friends frame than the Math frame (hypothesized to result from the greater prevalence of explicitly mathematical reasoning problems in the models’ training). To start, we want to see if these patterns hold for the other RLLMs we test, and how their error rates compare to o1-mini and DeepSeek-R1.
@@ 3.2 Methods
We test Claude 3.7 Sonnet, o3-mini (low, medium, and high), Gemini 2.5 Pro Preview, and Grok 3 Mini Beta (low and high) on Heyman and Zylberberg (2025)’s 4-vertex 2-coloring (“4v2c") and 8-vertex 4-coloring (“8v4c") problem sets. The 4v2c problem set consists of all possible 4-vertex graphs counting vertex-order permutations separately, minus the graph with no edges, for a total of $632^{C(4,2)}-1=$ 63 problems (40 colorable and 23 uncolorable). The version of the 8v4c problem set on which Heyman and Zylberberg (2025) test their RLLMs, which is also the version we use, consists of random samples of 50 8-vertex graphs from each possible edge count between 14 and 23 inclusive (out of a maximum of $C(8,2)=28$ edges), for a total of 500 problems (306 colorable and 194 uncolorable). Like Heyman and Zylberberg (2025), we prompt each model with each problem in each frame five separate times to increase statistical power given model stochasticity. We then evaluate the models’ responses and calculate the relevant error rates. We include the responses from o1-mini and DeepSeek-R1 collected by Heyman and Zylberberg (2025) (available under the MIT License) in our analysis for comparison.
@@ 3.3 Results
Rates of different error types for each model and frame on 8v4c are plotted in Figure 2 (note the varying y-axis scales). Errors are categorized into false colorable (giving a coloring for an uncolorable problem), wrong coloring (giving an invalid coloring for a problem that has valid colorings), and false uncolorable (declaring a colorable problem to be uncolorable). Error rates vary widely between models here, with Claude 3.7 S reaching $>10\%$ error on Math colorable problems and $>20\%$ error on Friends colorable problems, while o3-mini (high) and Grok 3 MB (high) each make no errors at all on Math, and only 2 and 5 errors on Friends respectively (out of 2500 total trials and 1530 colorable trials). However, the pattern of most errors (both in absolute terms and relative to applicable trials) being false-uncolorable reappears in almost all of the newer RLLMs alongside o1-mini and DeepSeek-R1; the only exception is Grok 3 MB (high), with an error sample size of only 5 (2 false uncolorable and 3 wrong coloring). Not surprisingly, both o3-mini and Grok 3 MB exhibit drastically decreasing error rates as reasoning effort increases. In terms of the relationship between frames, o1-mini and DeepSeek-R1’s pattern of more frequent errors on Friends than on Math is shared by Claude 3.7 Sonnet and (possibly, considering the confidence intervals) Gemini 2.5 PP, o3-mini (medium and high), and Grok 3 MB (high); however, Grok 3 MB (low) and (possibly) o3-mini (low) exhibit lower error on Friends instead.
@@ 4.3 Results
Figure 4 shows, for each frame and model with false-uncolorable false edge data, the proportion of false-uncolorable responses that include edge hallucinations, and the proportion that furthermore would answer correctly if the hallucinated edge(s) actually existed (i.e. the graph really would be uncolorable then). A response falling into the “would be correct" category is evidence (though not proof) that its edge hallucinations are the cause of its incorrectness. Depending on the model and frame, the “would be correct" proportion of responses can be as many as 33 out of 39 (DeepSeek-R1 Math) or 15 out of 16 (Gemini 2.5 PP Friends), or as low as 17 out of 143 (Grok 3 MB (low) Math) – but it is at least significantly above 0 in all cases with data available. Notably, for all combinations of frame and model with exposed CoT, there were many responses that included edge hallucinations but did not give false-uncolorable answers; causes of this can include the graph being uncolorable to begin with, the added edges still leaving the graph colorable, or the model’s chain of thought recognizing and correcting its mistake before completing (which we have informally observed in a considerable number of responses, but which would be difficult to systematically screen for).
Also of note is that, for every model-frame combination with the relevant data, the distribution of which edges get hallucinated is distinctly non-uniform, even though the problem graphs were generated in a completely vertex-symmetric manner and thus the distribution of which edges do not exist is roughly uniform. The tendency is for edges involving the vertices with the highest indices to get hallucinated most; in the Math frame, with vertices numbered from 0-7 inclusive, the edge that a given model most frequently hallucinates is generally (6, 7), with (5, 6) and to a lesser extent (5, 7) also popular. Similarly, in the Friends frame, the most popular edge to hallucinate is (George, Heather), followed by (Fran, George) and (Fran, Heather). Confirming the causal significance of this striking cross-model pattern is part of the goal of the next experiment we recount.

### v3

@@ Abstract
Large language models have recently made great strides in reasoning task performance through chain-of-thought (CoT) strategies trained via reinforcement learning; however, these “reasoning large language models" (RLLMs) remain imperfect reasoners, and understanding the frequencies and causes of their failure modes is important for both users and developers. We test o1-mini, o3-mini, DeepSeek-R1, Claude 3.7 Sonnet, Gemini 2.5 Pro Preview, and Grok 3 Mini Beta on graph coloring as a variable-complexity constraint-satisfaction logic problem, and find evidence from both error rate comparisons and CoT/explanation text analysis that RLLMs are prone to hallucinate graph edges not specified in the prompt. This phenomenon persists across multiple problem complexity levels and semantic frames, and it appears to account for a significant fraction of the incorrect answers from every tested model, and the vast majority of them for some models. We also validate the generalizability of this input-conflicting hallucination phenomenon with smaller-scale experiments on a type of stable matching problem. Our results indicate that RLLMs may possess broader issues with misrepresentation of problem specifics, and we offer suggestions for design choices to mitigate this weakness.
@@ 1 Introduction
The precise nature, capabilities, and limitations of RLLM reasoning is a subject of ongoing research, though it faces challenges both from the typical inscrutability of large neural networks and from the frequency with which developers keep model weights and/or training procedure details secret (Zhang et al., 2025a); some model APIs even leave the chain of thought completely internal and only report the model’s final answer, likely to prevent competitors from using the chains of thought as training data for their own models. Research so far has found that despite RLLMs’ improvements over traditional LLMs, they are still not entirely free of flaws such as hallucinating facts in their output (OpenAI, 2025c), reduced performance on problems not specifically included in the training corpus (Mirzadeh et al., 2024), reduced performance when superficially relevant distractor information is added to the prompt (Mirzadeh et al., 2024), imperfect performance even on problems that are simple and algorithmically solvable without fail (Valmeekam et al., 2024; Heyman and Zylberberg, 2025; Shojaee et al., 2025), decreasing performance with increasing problem complexity even as perfect accuracy remains algorithmically possible (Lin et al., 2025; Heyman and Zylberberg, 2025), and failure to correctly execute algorithms that are explicitly specified in the prompt (Shojaee et al., 2025). Furthermore, an RLLM’s chain of thought is not guaranteed to accurately report the process by which it arrives at its final answer (Chen et al., 2025b). Gathering more and better evidence about the frequencies and causes of RLLM reasoning failure modes is important both for users deploying existing models and for developers seeking to build more capable and reliable models.
@@
Figure 1: Left: A simple undirected graph with 4 vertices and 4 edges, shown with a valid 2-coloring. Right: A different 4-vertex, 4-edge graph that cannot be 2-colored.
@@
Some prior work (Stechly et al., 2024; Heyman and Zylberberg, 2025) has evaluated LLM reasoning through graph coloring problems, in which the model must attempt to assign one of $k$ colors to each of the $n$ vertices of a specified undirected graph such that no two vertices with an edge between them receive the same color. (See Figure 1 for a simple example.) What colorings (if any) are valid for a given graph depends on the graph’s precise arrangement of edges. A valid coloring (or proof that there is none) for a given graph and number of available colors $k$ can be found algorithmically without fail, though for $k\geq 3$ the problem is NP-complete.
@@
In our own prior work detailed in the preprint Heyman and Zylberberg (2025), we took interest in graph coloring as a logic problem type that tests both the ability to reason step-by-step (“if I assign these colors to these vertices, which colors can I then assign to which other vertices?") and the ability to systematically explore a space of possibilities (“if assigning these colors to these vertices forces a constraint violation, what other possible color assignments can I try, or have I already tried them all?"), and for which problem instances of widely varying complexity can be easily procedurally generated. While Stechly et al. (2024) test only the traditional LLM GPT-4 on graph coloring (as well as two other problem types), we tested the RLLMs o1-mini and DeepSeek-R1 (DeepSeek-AI, 2025a) alongside GPT-4o and three other traditional LLMs. We found that while the RLLMs perform much better than the traditional LLMs, the RLLMs remain imperfect even on simple 4-vertex 2-coloring problems, and their error rates increase dramatically as problem complexity increases to 8-vertex 4-coloring (the highest level we tested). In particular, at the 8-vertex level, both RLLMs’ errors mostly consist of claiming that graphs are uncolorable under the constraints when in reality valid colorings exist (as opposed to providing colorings that are in fact invalid, either on the dataset’s uncolorable problems or on its colorable ones). We also found a strong positive correlation between error rate and a measure of the difficulty of narrowing down a valid solution to a colorable problem, and we initially took this as evidence that possibility space exploration is RLLMs’ primary area of weakness in the context of problems like graph coloring. However, later validation experiments that we conducted yielded evidence that this correlation is not causally significant (see Appendix A for details), and thus the true primary causes of these RLLM errors likely lie elsewhere.
@@
In this work, we investigate the failure modes of RLLMs on graph coloring problems using not only comparison of error rates and types in the models’ final answers as in Heyman and Zylberberg (2025), but also analysis of the justifications they provide for those answers and of their chains of thought when available. Our experiments include o1-mini and DeepSeek-R1 as well as several newer RLLMs with superior overall benchmark performance. Remarkably, we find that all of the RLLMs we test are prone to hallucinate edges in the graph to be colored that are not listed in the prompt, which can lead them to mistake a colorable graph for an uncolorable one. (To use the terminology of Zhang et al. (2025b), these are input-conflicting hallucinations, reflecting a failure of models to properly deploy information they are explicitly given at runtime, which are distinct from the better-studied fact-conflicting hallucinations that reflect failure to properly memorize information during training.) This phenomenon persists across multiple problem complexity levels and multiple semantic frames in which graph coloring problems can be presented, and it appears to account for at least a significant fraction of the incorrect answers given by every model, and the vast majority of incorrect answers for some models. We present our findings across a series of experiments (including smaller-scale experiments that demonstrate a similar hallucination phenomenon in a type of stable matching problem; see Appendix D) and conclude by discussing our results’ broader implications for RLLM reasoning and how its flaws might be addressed.
@@ 2 General Methodology
We conduct our experiments on the following RLLMs: (1) OpenAI’s o1-mini (OpenAI, 2024b), specifically o1-mini-2024-09-12, through the OpenAI API; (2) OpenAI’s o3-mini (OpenAI, 2025b), specifically o3-mini-2025-01-31, through the OpenAI API; (3) DeepSeek-R1 (DeepSeek-AI, 2025a) through a combination of the Fireworks AI and DeepInfra online hosts; (4) Anthropic’s Claude 3.7 Sonnet (Anthropic, 2025), specifically claude-3-7-sonnet-20250219 with thinking mode turned on, through the Anthropic API; (5) Google’s Gemini 2.5 Pro Preview (Kavukcuoglu, 2025) through the Gemini Developer API; and (6) xAI’s Grok 3 Mini Beta (xAI, 2025) through the xAI API. For o3-mini and Grok 3 Mini Beta, we test all reasoning effort settings (low, medium, and high for the former, and low and high for the latter). We use the mini counterparts of o1, o3, and Grok 3 Beta rather than the full-size models for cost-effectiveness reasons and because the mini models are specialized for logical and mathematical reasoning; we do not expect the broader knowledge of the full-size models to be particularly helpful for the graph coloring task.
As in Heyman and Zylberberg (2025), for models with variable temperature settings, we set temperature to 0 to match the low-creativity deductive nature of the graph coloring task, with the exception of DeepSeek-R1; its developers recommend a temperature no lower than 0.5 (DeepSeek-AI, 2025b), and we use a temperature of 0.5 for this reason. We find non-terminating output from our prompts at these temperatures to be infrequent for DeepSeek-R1 and nonexistent for all other models.
@@
We use the same prompting format as in Heyman and Zylberberg (2025); the model is given a description of a graph, including a fully enumerated and clearly labeled set of edges, then asked if it is possible to color the graph with a specified set of colors so that no two adjacent vertices receive the same color, and to give a plan for it if so. To broaden our observations, we deploy prompts separately with two different frames, “Math" and “Friends", which we designed to probe LLM semantic biases. “Math" prompts present the problem explicitly as coloring a graph with edges connecting numbered vertices; meanwhile, in deliberately less-orthodox “Friends" prompts, the model is given a list of friendships that exist between $n$ people (named alphabetically as Alice, Bob, Carol, etc.) and told that everyone is going to wear a shirt with one of $k$ colors and no one wants to wear the same color shirt as any of their friends. (Examples of the prompts we use can be found in Appendix B.) When evaluating model responses for answer correctness, we use an automatic parser and manually review and interpret any responses in which it does not detect a coherent answer.
@@
Code for reproducing our experiments and records of our problems, prompts, and responses (including those in appendices) can be found at https://github.com/AlexHeyman/RLLMGraphColoring. The vast majority of the compute used in our experiments was on the model API side, making estimation of its magnitude difficult. However, each experiment section specifies the number of prompt submissions to the RLLMs involved; the totals across all experiments are 5,155 for o1-mini, 7,155 for DeepSeek-R1, 7,651 for Claude 3.7 S, 9,230 for Grok 3 MB (low), 14,230 for Grok 3 MB (high), 12,230 for o3-mini (high), 7,230 for each of o3-mini (low), o3-mini (medium), and Gemini 2.5 PP, and 2,000 for each of gpt-oss-120b (low), (medium), and (high) in Appendix D. All of our data analysis takes no more than 1 hour to run on a single AMD Ryzen 9 8945HS CPU.
@@ 3.1 Purpose
In Heyman and Zylberberg (2025), we find several basic patterns in the errors made by both o1-mini and DeepSeek-R1, including (1) errors on 4-vertex 2-coloring problems are very rare but not impossible; (2) on 8-vertex 4-coloring problems, errors are much more frequent and mostly consist of declaring colorable problems to be uncolorable; and (3) error rates are generally higher for the Friends frame than the Math frame (which we hypothesize results from the greater prevalence of explicitly mathematical reasoning problems in the models’ training). To start, we want to see if these patterns hold for the other RLLMs we test in this work, and how their error rates compare to o1-mini and DeepSeek-R1.
@@ 3.2 Methods
We test Claude 3.7 Sonnet, o3-mini (low, medium, and high), Gemini 2.5 Pro Preview, and Grok 3 Mini Beta (low and high) on the 4-vertex 2-coloring (“4v2c") and 8-vertex 4-coloring (“8v4c") problem sets on which we also tested o1-mini and DeepSeek-R1 in Heyman and Zylberberg (2025). The 4v2c problem set consists of all possible 4-vertex graphs counting vertex-order permutations separately, minus the graph with no edges, for a total of $2^{C(4,2)}-1=63$ problems (40 colorable and 23 uncolorable). The version of the 8v4c problem set on which we test our RLLMs consists of random samples of 50 8-vertex graphs from each possible edge count between 14 and 23 inclusive (out of a maximum of $C(8,2)=28$ edges), for a total of 500 problems (306 colorable and 194 uncolorable). We prompt each model with each problem in each frame five separate times to increase statistical power given model stochasticity. We then evaluate the models’ responses and calculate the relevant error rates. We include the responses from o1-mini and DeepSeek-R1 that we analyzed in Heyman and Zylberberg (2025) in our analysis here as well.
@@ 3.3 Results
Rates of different error types for each model and frame on 8v4c are plotted in Figure 2 (note the varying y-axis scales). Errors are categorized into false colorable (giving a coloring for an uncolorable problem), wrong coloring (giving an invalid coloring for a problem that has valid colorings), and false uncolorable (declaring a colorable problem to be uncolorable). Error rates vary widely between models here, with Claude 3.7 S reaching $10\%$ error on Math colorable problems and $20\%$ error on Friends colorable problems, while o3-mini (high) and Grok 3 MB (high) each make no errors at all on Math, and only 2 and 5 errors on Friends respectively (out of 2500 total trials and 1530 colorable trials). However, the pattern of most errors (both in absolute terms and relative to applicable trials) being false-uncolorable reappears in almost all of the newer RLLMs alongside o1-mini and DeepSeek-R1; the only exception is Grok 3 MB (high), with an error sample size of only 5 (2 false uncolorable and 3 wrong coloring). Not surprisingly, both o3-mini and Grok 3 MB exhibit drastically decreasing error rates as reasoning effort increases. In terms of the relationship between frames, o1-mini and DeepSeek-R1’s pattern of more frequent errors on Friends than on Math is shared by Claude 3.7 Sonnet and (possibly, considering the confidence intervals) Gemini 2.5 PP, o3-mini (medium and high), and Grok 3 MB (high); however, Grok 3 MB (low) and (possibly) o3-mini (low) exhibit lower error on Friends instead.
@@ 4.3 Results
Figure 4 shows, for each frame and model with false-uncolorable false edge data, the proportion of false-uncolorable responses that include edge hallucinations, and the proportion that furthermore would answer correctly if the hallucinated edge(s) actually existed (i.e. the graph really would be uncolorable then). A response falling into the “would be correct" category is evidence (though not proof) that its edge hallucinations are the cause of its incorrectness. Depending on the model and frame, the “would be correct" proportion of false-uncolorable responses can be as many as 33 out of 39 (DeepSeek-R1 Math) or 15 out of 16 (Gemini 2.5 PP Friends), or as low as 17 out of 143 (Grok 3 MB (low) Math) – but it is at least significantly above 0 in all cases with data available. Notably, for all combinations of frame and model with exposed CoT, there were many responses that included edge hallucinations but did not give false-uncolorable answers; causes of this can include the graph being uncolorable to begin with, the added edges still leaving the graph colorable, or the model’s chain of thought recognizing and correcting its mistake before completing (which we have informally observed in a considerable number of responses, but which would be difficult to systematically screen for).
Also of note is that, for every model-frame combination with the relevant data available, the distribution of which edges get hallucinated is distinctly non-uniform, even though the problem graphs were generated in a completely vertex-symmetric manner and thus the distribution of which edges do not exist is roughly uniform. The tendency is for edges involving the vertices with the highest indices to get hallucinated most; in the Math frame, with vertices numbered from 0-7 inclusive, the edge that a given model most frequently hallucinates is generally (6, 7), with (5, 6) and to a lesser extent (5, 7) also popular. Similarly, in the Friends frame, the most popular edge to hallucinate is (George, Heather), followed by (Fran, George) and (Fran, Heather). Confirming the causal significance of this striking cross-model pattern is part of the goal of the next experiment we recount.